OpenStudy (anonymous):

WIN MEDALLLL!!!! NEED HELP!!! Which three ordered pairs complete the table to give solutions of the equation y = x – 7? Options- A. (–3, –10), (0, –7), and (3, 4) B. (–3, –10), (0, –7), and (3, –4) C. (–3, –4), (0, –7), and (3, –4) D. (–3, 4), (0, –7), and (3, –4)

OpenStudy (anonymous):

HELP!!

OpenStudy (anonymous):

Ok. Let's choose x = 1, what is y ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Can you give me the solution to the answer????

OpenStudy (anonymous):

Nope, you should solve it, I can help you.

OpenStudy (anonymous):

I dont get it!! Thats why!

OpenStudy (anonymous):

Ok, the eqn is y = x -7. right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now put x= 1 in the eqn you get y = 1-7

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Tell me what is y

OpenStudy (anonymous):

1-7?

OpenStudy (anonymous):

1- 7 = -6

OpenStudy (anonymous):

Choose some other values of x and find y

OpenStudy (anonymous):

Your first ordered pair is (1,-6)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

But my options are- A. (–3, –10), (0, –7), and (3, 4) B. (–3, –10), (0, –7), and (3, –4) C. (–3, –4), (0, –7), and (3, –4) D. (–3, 4), (0, –7), and (3, –4)

OpenStudy (anonymous):

You have 3 ordered pairs in each solution. You have to find the option in which the 3 ordered pairs match the equation. Remember that an ordered pair is in the form (x , y) where x is the first term, and then y is after the comma. therefore you take each ordered pair, starting with the first one and we have (-3, -10) so x = -3 y = -10 Now if we plug these values into the equation, does it prove true? y = x -7 -10 = -3 - 7 so this is true, so you have your first ordered pair checks out. Now you have to check the rest of them, until you find the option where all 3 are correct.

OpenStudy (anonymous):

I'm pretty sure its B. I graphed it.

OpenStudy (anonymous):

Yes, it's B. Good Job ! @lizziebeth